Russell's has annual revenue of $387,000 with costs of $216,400. Depreciation is $48,900 and the tax rate is 21 percent. The firm has debt outstanding with a market value of $182,000 along with 9,500 shares of stock that is selling at $67 a share. The firm has $48,000 of cash of which $29,500 is needed to run the business. 


What is the firm's EV/EBITDA ratio?
Answer #1
Enterprise value = Market value of debt + Market value of equity - Cash = 182000 + (9500*67) - 48000 770500
EBITDA = Annual revenue - Costs = 387000 - 216400 170600
EV/EBITDA ratio = 770500 / 170600 4.52